P(3,-3) lies on the terminal arm of an angle in standard position.

a. Draw a sketch of the principal angle (θ) in standard position. Label θ AND the related acute

angle, RA. [2 marks]

b. Determine the exact primary trig ratios for θ. Show ALL your work. You do not need to

rationalize the denominator, but you must simplify your radical and your ratios. [5 marks]

c. Determine the values of the principal angle and related acute angle to the nearest degree.


1
Expert's answer
2022-01-10T16:37:49-0500

A)








B)

"y=\sqrt{r^2-x^2}=\sqrt{7^2-2^2}=\sqrt{49-4}=\sqrt{45}=3\sqrt{5}"

"sin \theta=-\frac{y}{r}=-\frac{3 \sqrt{5}}{7}"

"tan \theta=\frac{y}{x}=\frac{3 \sqrt{5}}{2}"

c) "tan \theta=\frac{y}{x}=\frac{3 \sqrt{5}}{2} \implies \theta= tan^{-1}\frac{3 \sqrt{5}}{2}=73.4"

"\theta_{Total}=180+73.4=253.4 =253^0"